# Preasymptotic error estimates of higher-order EEM for the time-harmonic Maxwell equations with large wave number

### v1

@@ 4.1. $\kappa$-explicit regularity by decomposition for the Maxwell dual problem
Suppose that $\Omega$ is $C^{2}$. There exists a constant $\tilde{C}_{0}>0$ such that if $\kappa h\leq\tilde{C}_{0}$, then the following error estimates hold for any ${\bm{u}}\in\bm{V}$ and $\bm{v}_{h}\in\bm{V}_{h}$,
@@
$b({\bm{w}},{\bm{\phi}})=\left(P^{\perp}\bm{v},{\bm{\phi}}\right)\quad\forall{\bm{\phi}}\in\bm{V},$
@@
This superconvergent order error estimate in the negative norm are quite similar to the standard negative norm error estimate of higher order FEM for the elliptic problems (see e.g. [8]).
@@ 4.2. Stability estimates of EEM
and by Theorem 2.1, Lemmas 4.14.2 and (4.20), there hold the following estimates:
@@
$\|({\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}})_{T}\|_{\Gamma}$ $\lesssim h^{\frac{3}{2}}\big{(}\|{\bm{w}}^{0}\|+\kappa^{-1}\|{\bm{w}}^{0}\|_{1}\big{)}+\kappa^{p}h^{p+\frac{1}{2}}\|{\bm{w}}^{0}\|,$ (4.27)
$\inf_{\bm{v}_{h}\in\bm{V}_{h}}\interleave{\bm{\Psi}}-\bm{v}_{h}\interleave$ $\lesssim h\|{\bm{w}}^{0}\|+\kappa^{-1}h\|{\bm{w}}^{0}\|_{1}+\kappa^{p}h^{p}\|{\bm{w}}^{0}\|$
$\lesssim(h+\kappa^{p}h^{p})\|{\bm{w}}^{0}\|+\kappa^{-1}\|{\bm{E}}_{h}\|,$ (4.28)
since $\kappa h\leq\tilde{C}_{0}$ can be guaranteed if $\kappa^{2p+1}h^{2p}$ is sufficiently small. Using (4.24)–(4.28) and Lemmas 3.1, 4.2 and 4.3, we deduce that
@@
$\lesssim\|\bm{f}\|\big{(}\|{\bm{\Psi}}\|+h^{2}\|{\bm{\Psi}}\|_{2}\big{)}+\|\bm{g}\|_{\Gamma}\big{(}\|{\bm{\Psi}}_{T}\|_{\Gamma}+\|({\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}})_{T}\|_{\Gamma}\big{)}$
@@
$\lesssim\kappa^{-1}\|\bm{f}\|\|{\bm{w}}^{0}\|+\kappa h^{2}\|\bm{f}\|\|{\bm{w}}^{0}\|+\kappa^{-1}h\|\bm{f}\|\|{\bm{E}}_{h}\|+\kappa^{-1}\|\bm{g}\|_{\Gamma}\|{\bm{w}}^{0}\|$
$\quad+\|\bm{g}\|_{\Gamma}\big{(}h^{\frac{3}{2}}\|{\bm{w}}^{0}\|+\kappa^{-1}h^{\frac{3}{2}}\|{\bm{w}}^{0}\|_{1}+\kappa^{p}h^{p+\frac{1}{2}}\|{\bm{w}}^{0}\|\big{)}+\kappa^{p+1}h^{p}\inf_{\bm{v}_{h}\in\bm{V}_{h}}\interleave{\bm{\Psi}}-\bm{v}_{h}\interleave\|\bm{E}_{h}\|$
$\lesssim\kappa^{-1}\|\bm{f}\|\|{\bm{w}}^{0}\|+\kappa^{-1}h\|\bm{f}\|\|{\bm{E}}_{h}\|+\kappa^{-1}\|\bm{g}\|_{\Gamma}\|{\bm{w}}^{0}\|+\kappa^{-1}h^{\frac{1}{2}}\|\bm{g}\|_{\Gamma}\|{\bm{E}}_{h}\|$
$\quad+\kappa^{p+1}h^{p}\|\bm{E}_{h}\|\big{(}(h+\kappa^{p}h^{p})\|{\bm{w}}^{0}\|+\kappa^{-1}\|{\bm{E}}_{h}\|\big{)}$
$\lesssim\kappa^{-1}\big{(}\|\bm{f}\|+\|\bm{g}\|_{\Gamma}\big{)}\|{\bm{w}}^{0}\|+(\kappa^{2p+1}h^{2p}+\kappa^{p+1}h^{p+1})\|{\bm{w}}^{0}\|\|{\bm{E}}_{h}\|+\kappa^{p}h^{p}\|{\bm{E}}_{h}\|^{2}$
@@
$\|{\bm{w}}^{0}\|^{2}$ $\lesssim\|{\bm{w}}^{0}-{\bm{w}}_{h}^{0}\|^{2}+\|{\bm{E}}_{h}^{0}-{\bm{E}}_{h}\|^{2}+\kappa^{-1}\big{(}\|\bm{f}\|+\|\bm{g}\|_{\Gamma}\big{)}\|{\bm{w}}^{0}\|+(\kappa^{2p+1}h^{2p}+\kappa^{p+1}h^{p+1})\|\bm{E}_{h}\|\|{\bm{w}}^{0}\|$
@@ 4.3. Preasymptotic error estimates of EEM
Without loss of generality, suppose $\|\mathcal{T}\textsf{P}^{\perp}{\bm{\eta}}\|\neq 0$. From Lemmas 3.1 and 4.3, we have
@@
  • It reveals the advantages of higher-order EEM applied to the high frequency Maxwell problem in the preasymptotic regime. In contrast, the theories presented in [42, 43] primarily focus on the quasi-optimality of $hp$-EEM, i.e., in the asymptotic regime.
@@
(c) For problems with large wave number, a discrete solution of reasonable accuracy requires the pollution error $\kappa^{2p+1}h^{2p}$ to be small enough (see, e.g. [2, 4, 31]). From this point of view, our mesh condition $\kappa^{2p+1}h^{2p}\leq C_{0}$ is quite practical.
(d) The tailor-made truncation operator $\mathcal{T}$ is crucial for weakening the mesh condition (see Lemma 4.3). If the elliptic projections were replaced by those given in [38, (3.8)-(3.9)] instead of (3.6) and (4.9), the best we could achieve would be a preasymptotic error estimate under the condition that $\kappa^{p+2}h^{p+1}$ is sufficiently small.

### v2

@@ 3.3. Elliptic projections
Suppose that $\Omega$ is $C^{2}$. There exists a constant $\tilde{C}_{0}>0$ such that if $\kappa h\leq\tilde{C}_{0}$, then the following error estimates hold for any ${\bm{u}}\in\bm{V}$ and $\bm{v}_{h}\in\bm{V}_{h}$:
@@
$b({\bm{w}},{\bm{\phi}})=\left(\textsf{P}^{\perp}\bm{v},{\bm{\phi}}\right)\quad\forall{\bm{\phi}}\in\bm{V},$
@@
This error estimate in the negative norm is quite similar to the standard negative norm error estimate of higher-order FEM for the elliptic problems (see e.g. [8]).
4. Preasymptotic error estimates for the EEM
In this section, we first establish the $\kappa$-explicit regularity of the solution to the Maxwell problem (1.1) by decomposing it into a non-oscillatory elliptic part and an oscillatory yet smooth part. Based on this decomposition, we then establish wave-number-explicit stability estimates for the $p$-th order EEM (2.5) using the “modified duality argument” [58]. Finally, we use these stability estimates to derive preasymptotic error estimates for the EEM.
4.1. $\kappa$-explicit regularity by decomposition for the Maxwell problem
The following theorem says that the solution $\bm{E}$ to the Maxwell problem (1.1) can be decomposed into the sum of an elliptic part and a smooth part: $\bm{E}={\bm{E}}_{\mathcal{E}}+{\bm{E}}_{\mathcal{A}}$, where $\bm{E}$$_{\mathcal{E}}$ is non-oscillatory in the sense that its ${\bm{H}}^{2}$-norm is uniformly bounded for all $\kappa$} and ${\bm{E}}_{\mathcal{A}}$ is oscillatory but the ${\bm{H}}^{j}$-bound of ${\bm{E}}_{\mathcal{A}}$ is available for any integer $j\geq 0$ when the domain is sufficiently smooth.
Theorem 4.1 (Regularity decomposition).
Suppose that $\Omega$ is $C^{m}$ for some integer $m\geq 2$ and $\bm{f}\in{\bm{H}}(\operatorname{{\mathrm{div}}}^{0};\Omega)$, $\bm{g}=\bm{g}_{T}\in{\bm{H}}^{\frac{1}{2}}(\Gamma)$. Let $\bm{E}$ be the solution to problem (1.1), then there exists a splitting $\bm{E}={\bm{E}}_{\mathcal{E}}+{\bm{E}}_{\mathcal{A}}$ such that ${\bm{E}}_{\mathcal{E}}\in{\bm{H}}^{1}(\Omega)$ and ${\bm{E}}_{\mathcal{A}}\in{\bm{H}}^{m}(\Omega)$ satisfying
$\kappa^{1-m}\|{\bm{E}}_{\mathcal{A}}\|_{m}+\kappa^{2}\|{\bm{E}}_{\mathcal{E}}\|+\kappa\|{\bm{E}}_{\mathcal{E}}\|_{1}\lesssim\|\bm{f}\|+\|\bm{g}\|_{\frac{1}{2},\Gamma}.$ (4.1)
Furthermore, if $\bm{f}\cdot\bm{\nu}\in{\bm{H}}^{\frac{1}{2}}(\Gamma)$ and $\operatorname{{\mathrm{div}}}_{\Gamma}\bm{g}\in{\bm{H}}^{\frac{1}{2}}(\Gamma)$, then ${\bm{E}}_{\mathcal{E}}\in{\bm{H}}^{2}(\Omega)$ and
$\|{\bm{E}}_{\mathcal{E}}\|_{2}\lesssim\|\bm{f}\|+\kappa^{-1}\|\bm{f}\cdot\bm{\nu}\|_{\frac{1}{2},\Gamma}+\|\bm{g}\|_{\frac{1}{2},\Gamma}+\kappa^{-1}\|\operatorname{{\mathrm{div}}}_{\Gamma}\bm{g}\|_{\frac{1}{2},\Gamma}.$ (4.2)
Proof.
We begin by introducing a boundary frequency splitting of $\bm{g}$, inspired by the construction in [43, Section 6]. Let $\bm{g}=\bm{g}_{T}=(g_{1},g_{2},g_{3})^{T}\in{\bm{H}}^{\frac{1}{2}}(\Gamma)$. For each component $g_{i}\in H^{1/2}(\Gamma)$ $(i=1,2,3)$, we define its harmonic extension $\mathcal{E}_{\Omega}g_{i}\in H^{1}(\Omega)$ as the weak solution of
$\begin{cases}\Delta(\mathcal{E}_{\Omega}g_{i})=0,&\text{in }\Omega,\\ \mathcal{E}_{\Omega}g_{i}|_{\Gamma}=g_{i},&\text{on }\Gamma.\end{cases}$ (4.3)
By elliptic regularity (cf. [39]), the operator $\mathcal{E}_{\Omega}$ extends continuously as $\mathcal{E}_{\Omega}:H^{s-\frac{1}{2}}(\Gamma)\to H^{s}(\Omega),\;1\leq s\leq m,$ and hence
$\|\mathcal{E}_{\Omega}g_{i}\|_{s}\lesssim\|g_{i}\|_{s-\frac{1}{2},\Gamma},\qquad 1\leq s\leq m.$ (4.4)
Next, following [41, Section 4], using the frequency filters $H_{\Omega}$ and $L_{\Omega}$ there, we have the splitting
$\mathcal{E}_{\Omega}g_{i}=H_{\Omega}\mathcal{E}_{\Omega}g_{i}+L_{\Omega}\mathcal{E}_{\Omega}g_{i},\qquad i=1,2,3.$ (4.5)
By Lemmas 4.2–4.3 of [41], the two components satisfy the bounds
$\|H_{\Omega}\mathcal{E}_{\Omega}g_{i}\|_{s^{\prime}}$ $\lesssim\kappa^{s^{\prime}-s}\|\mathcal{E}_{\Omega}g_{i}\|_{s},$  $\forall\,0\leq s^{\prime}\leq s,$ (4.6)
$\|L_{\Omega}\mathcal{E}_{\Omega}g_{i}\|_{s}$ $\lesssim\kappa^{s-1}\|\mathcal{E}_{\Omega}g_{i}\|_{1},$ $\forall\,s\geq 1.$ (4.7)
Define
${\bm{G}}_{\mathcal{E}}=(H_{\Omega}\mathcal{E}_{\Omega}g_{1},H_{\Omega}\mathcal{E}_{\Omega}g_{2},H_{\Omega}\mathcal{E}_{\Omega}g_{3})^{T},\qquad{\bm{G}}_{\mathcal{A}}=(L_{\Omega}\mathcal{E}_{\Omega}g_{1},L_{\Omega}\mathcal{E}_{\Omega}g_{2},L_{\Omega}\mathcal{E}_{\Omega}g_{3})^{T}.$
We set ${\bm{g}}_{\mathcal{E}}=({\bm{G}}_{\mathcal{E}})_{T}$ and ${\bm{g}}_{\mathcal{A}}=({\bm{G}}_{\mathcal{A}})_{T}$. Clearly, $\bm{g}={\bm{g}}_{\mathcal{E}}+{\bm{g}}_{\mathcal{A}}$ on $\Gamma$.
By the trace inequality together with (4.6)–(4.7) and (4.4), we obtain
$\|{\bm{g}}_{\mathcal{E}}\|_{\Gamma}\lesssim\|{\bm{G}}_{\mathcal{E}}\|_{1}^{1/2}\|{\bm{G}}_{\mathcal{E}}\|^{1/2}{\lesssim\kappa^{-\frac{1}{2}}\sum_{i=1}^{3}\left\|\mathcal{E}_{\Omega}g_{i}\right\|_{1}}\lesssim\kappa^{-\frac{1}{2}}\sum_{i=1}^{3}\left\|g_{i}\right\|_{\frac{1}{2},\Gamma}\lesssim\kappa^{-\frac{1}{2}}\|\bm{g}\|_{\frac{1}{2},\Gamma};$ (4.8)
$\|{\bm{g}}_{\mathcal{E}}\|_{j+\frac{1}{2},\Gamma}\lesssim\|{\bm{G}}_{\mathcal{E}}\|_{j+1}\lesssim\sum_{i=1}^{3}\left\|\mathcal{E}_{\Omega}g_{i}\right\|_{j+1}\lesssim\|\bm{g}\|_{j+\frac{1}{2},\Gamma},\qquad 0\leq j\leq 1;$ (4.9)
$\|{\bm{g}}_{\mathcal{A}}\|_{j+\frac{1}{2},\Gamma}\lesssim\|{\bm{G}}_{\mathcal{A}}\|_{j+1}\lesssim\kappa^{j}\sum_{i=1}^{3}\left\|\mathcal{E}_{\Omega}g_{i}\right\|_{1}\lesssim\kappa^{j}\|\bm{g}\|_{\frac{1}{2},\Gamma},\qquad 0\leq j\leq m-1.$ (4.10)
A triangle inequality and (4.10) yields that
$\|\operatorname{{\mathrm{div}}}_{\Gamma}{\bm{g}}_{\mathcal{E}}\|_{\frac{1}{2},\Gamma}\lesssim\|\operatorname{{\mathrm{div}}}_{\Gamma}\bm{g}\|_{\frac{1}{2},\Gamma}+\|{\bm{g}}_{\mathcal{A}}\|_{\frac{3}{2},\Gamma}\lesssim\|\operatorname{{\mathrm{div}}}_{\Gamma}\bm{g}\|_{\frac{1}{2},\Gamma}+\kappa\|\bm{g}\|_{\frac{1}{2},\Gamma}.$ (4.11)
We now turn to the regularity decomposition of the Maxwell solution $\bm{E}$ by following the idea proposed in [34]. Let ${\bm{E}}_{\mathcal{E}}\in\bm{V}$ be the solution of
$b({\bm{E}}_{\mathcal{E}},\bm{v})=(\bm{f},\bm{v})+\langle{\bm{g}}_{\mathcal{E}},\bm{v}_{T}\rangle\qquad\forall\bm{v}\in\bm{V}.$
Then the estimates for $\bm{E}$$_{\mathcal{E}}$ in (4.1)–(4.2) follow from Lemma 3.3 together with (4.8)–(4.9) and (4.11).
Define ${\bm{E}}_{\mathcal{A}}=\bm{E}-{\bm{E}}_{\mathcal{E}}$. It follows that ${\bm{E}}_{\mathcal{A}}$ solves
$a({\bm{E}}_{\mathcal{A}},\bm{v})=C_{\mathcal{T}}\kappa^{2}\big(\mathcal{T}\textsf{P}^{\perp}{\bm{E}}_{\mathcal{E}},\bm{v}\big)+\langle{\bm{g}}_{\mathcal{A}},\bm{v}_{T}\rangle,\qquad\forall\bm{v}\in\bm{V}.$
By Theorem 2.1, the stability of $\mathcal{T}$ and $\textsf{P}^{\perp}$, the estimate of $\bm{E}$$_{\mathcal{E}}$ and (4.10), we deduce
$\kappa^{-1}\|{\bm{E}}_{\mathcal{A}}\|_{1}+\|{\bm{E}}_{\mathcal{A}}\|$ $\lesssim C_{\mathcal{T}}\kappa\|\mathcal{T}\textsf{P}^{\perp}{\bm{E}}_{\mathcal{E}}\|+\kappa^{-1}\|{\bm{g}}_{\mathcal{A}}\|_{\Gamma}+\kappa^{-2}\|{\bm{g}}_{\mathcal{A}}\|_{\frac{1}{2},\Gamma}\lesssim\kappa^{-1}(\|\bm{f}\|+\|\bm{g}\|_{\frac{1}{2},\Gamma}),$
$\|{\bm{E}}_{\mathcal{A}}\|_{m}$ $\lesssim C_{\mathcal{T}}\kappa^{m+1}\Big(\|\mathcal{T}\textsf{P}^{\perp}{\bm{E}}_{\mathcal{E}}\|+\kappa^{-m+1}\|\mathcal{T}\textsf{P}^{\perp}{\bm{E}}_{\mathcal{E}}\|_{m-2,\kappa}+\kappa^{-m}\|(\mathcal{T}\textsf{P}^{\perp}{\bm{E}}_{\mathcal{E}})\cdot\bm{\nu}\|_{m-\frac{3}{2},\Gamma}\Big)$
$\quad+\kappa^{m-1}\|{\bm{g}}_{\mathcal{A}}\|_{\Gamma}+\|{\bm{g}}_{\mathcal{A}}\|_{m-\frac{3}{2},\kappa,\Gamma}+\kappa^{-1}\|\operatorname{{\mathrm{div}}}_{\Gamma}{\bm{g}}_{\mathcal{A}}\|_{m-\frac{3}{2},\kappa,\Gamma}$
$\lesssim\kappa^{m-1}\big(\|\bm{f}\|+\|\bm{g}\|_{\frac{1}{2},\Gamma}\big).$ (4.12)
This completes the proof. ∎
Remark 4.1.
(a) As discussed in [24, 34], the above proof presents a more technically efficient method compared to existing regularity splitting approaches, such as the Fourier technique [40, 41, 42, 43] and iterative techniques found in [16, 17].
(b) Alternatively, one may employ the boundary frequency filters in [43, Eq. (6.16)] to obtain the same estimates but under an analytic boundary regularity assumption. The technique, however, does not extend to domains with merely $C^{m}$ boundaries, since the construction involves the harmonic extension of scalar surface potentials and analytic continuation techniques, which cannot be applied under our smoothness assumption.
(c) When $\operatorname{{\mathrm{div}}}\bm{f}\neq 0$, the estimates for $\|{\bm{E}}_{\mathcal{A}}\|_{m}$ and $\|{\bm{E}}_{\mathcal{E}}\|$ in (4.1) still hold. The estimates for $\kappa\|{\bm{E}}_{\mathcal{E}}\|_{1}$ in (4.1) still holds if $\kappa^{-1}\|\operatorname{{\mathrm{div}}}\bm{f}\|$ is added to the right-hand side of (4.1). And the estimate (4.2) remain valid if $\kappa^{-1}\|\operatorname{{\mathrm{div}}}\bm{f}\|+\kappa^{-2}\|\operatorname{{\mathrm{div}}}\bm{f}\|_{1}$ is added to its right-hand side and $\Omega$ is $C^{3}$ (see Remark 2.2 (b)).
@@ 4.2. Stability estimates of EEM
and by Theorems 2.1 and 4.1 and (4.19), there exists the decomposition ${\bm{\Psi}}={{\bm{\Psi}}}_{\mathcal{E}}+{{\bm{\Psi}}}_{\mathcal{A}}$, and the following estimates hold:
@@
${\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{2}}$ $\lesssim\|{\bm{w}}^{0}\|+\kappa^{-1}\|{\bm{w}}^{0}\|_{1}\lesssim\|{\bm{w}}^{0}\|+\kappa^{-1}h^{-1}\|{\bm{E}}_{h}\|,$ (4.26)
${\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{p+1}}$ $\lesssim\kappa^{p}\|{\bm{w}}^{0}\|.$ (4.27)
Since $\kappa h\leq\tilde{C}_{0}$ can be guaranteed if $\kappa^{2p+1}h^{2p}$ is sufficiently small, from Lemmas 2.1 and 3.43.5, and (4.26)–(4.27), we have
$\|{\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}}\|+h^{\frac{1}{2}}\|({\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}})_{T}\|_{\Gamma}+h^{1-p}\|\textsf{P}^{\perp}({\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}})\|_{-p+1,\kappa}$
$\lesssim\|{\bm{\Psi}}-{\bm{\Pi}_{h}}{\bm{\Psi}}\|+h^{\frac{1}{2}}\|({\bm{\Psi}}-{\bm{\Pi}_{h}}{\bm{\Psi}})_{T}\|_{\Gamma}+h\interleave{\bm{\Psi}}-{\bm{\Pi}_{h}}{\bm{\Psi}}\interleave$
$\lesssim h^{2}\|{{\bm{\Psi}}}_{\mathcal{E}}\|_{2}+h^{p+1}\|{{\bm{\Psi}}}_{\mathcal{A}}\|_{p+1}$
$\lesssim h\big((h+\kappa^{p}h^{p})\|{\bm{w}}^{0}\|+\kappa^{-1}\|{\bm{E}}_{h}\|\big).$ (4.28)
Using (4.23)–(4.25), (4.28), and Lemma 3.1, we deduce that
@@
$\lesssim{\|\bm{f}\|\|{\bm{\Psi}}\|+\|\bm{g}\|_{\Gamma}\|{\bm{\Psi}}_{T}\|_{\Gamma}+\|\bm{f}\|\|{\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}}\|+\|\bm{g}\|_{\Gamma}\|({\bm{\Psi}}-{\bm{P}^{-}_{h}\bm{\Psi}})_{T}\|_{\Gamma}}$
@@
$\lesssim\kappa^{-1}\big(\|\bm{f}\|+\|\bm{g}\|_{\Gamma}\big)\|{\bm{w}}^{0}\|$
$\quad+{\big((h+\kappa^{p}h^{p})\|{\bm{w}}^{0}\|+\kappa^{-1}\|{\bm{E}}_{h}\|\big)\big(h\|\bm{f}\|+h^{\frac{1}{2}}\|\bm{g}\|_{\Gamma}+\kappa^{p+1}h^{p}\|\bm{E}_{h}\|\big)}$
$\lesssim\kappa^{-1}\big(\|\bm{f}\|+\|\bm{g}\|_{\Gamma}\big)\|{\bm{w}}^{0}\|+(\kappa^{p+1}h^{p+1}+\kappa^{2p+1}h^{2p})\|{\bm{w}}^{0}\|\|{\bm{E}}_{h}\|+\kappa^{p}h^{p}\|{\bm{E}}_{h}\|^{2}$
@@
$\|{\bm{w}}^{0}\|^{2}$ $\lesssim\|{\bm{w}}^{0}-{\bm{w}}_{h}^{0}\|^{2}+\|{\bm{E}}_{h}^{0}-{\bm{E}}_{h}\|^{2}+\kappa^{-1}\big(\|\bm{f}\|+\|\bm{g}\|_{\Gamma}\big)\|{\bm{w}}^{0}\|+(\kappa^{p+1}h^{p+1}+\kappa^{2p+1}h^{2p})\|\bm{E}_{h}\|\|{\bm{w}}^{0}\|$
@@ 4.3. Preasymptotic error estimates of EEM
Without loss of generality, suppose $\|\mathcal{T}\textsf{P}^{\perp}{\bm{\eta}}\|\neq 0$. From Lemmas 3.1 and 3.5, we have
@@
  • It reveals the advantages of higher-order EEM applied to the high frequency Maxwell problem in the preasymptotic regime. In contrast, the theories presented in [42, 43] primarily focus on the quasi-optimality of hphp-EEM, i.e., in the asymptotic regime.
@@
(c) The tailor-made truncation operator $\mathcal{T}$ is crucial for weakening the mesh condition (see Lemma 3.5). If the elliptic projections were replaced by those given in [38, (3.8)-(3.9)] instead of (3.6) and (3.17), the best result we could currently achieve would be a preasymptotic error estimate, under the condition that $\kappa^{p+2}h^{p+1}$ is sufficiently small.
(d) For problems with large wave number, a discrete solution of reasonable accuracy requires the pollution error $\kappa^{2p+1}h^{2p}$ to be small enough (see, e.g. [2, 4, 31]). From this point of view, our mesh condition $\kappa^{2p+1}h^{2p}\leq C_{0}$ is quite practical.